6. If the mass of an object is quadrupled and the speed decreased by a factor of three, by what factor does the kinetic energy change?

(a) 4/3 (b) 16/3

(c) 4/9 (d) 16/9



1
Expert's answer
2022-04-17T17:08:27-0400

The kinetic energy is given as follows:


"K = \dfrac{mv^2}{2}"

New speed and mass are: "m_1 = 4m, v_1 = v\/3". Thus, the new kinetic energy is:


"K_1 = \dfrac{m_1v_1}{2} = \dfrac{4m(v/3)^2}{2} = \dfrac{4}{9}\cdot \dfrac{mv^2}{2} = \dfrac{4}{9}K"

Thus, the kinetic energy change by factor of 4/9.


Answer. (c) 4/9.